Calculate the potential of the indicator electrode when 40% of Fe2+ has been titrated with Ce(IV).

Fe2+ +Ce4+ <--> Fe3+ + Ce3+

(Fe3+/Fe2+): E=0.771 V

(Ce4+/Ce3+): E=1.44 V
Answer #1

Given +2 13 fe ce fel - + C 37 osthoorly (s} E celu z ÉCathoda - famode B = { cell/ (63+ - € +31 (a +2 = 1.44- 0.771 Ecen =